OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5

OpenStudy (anonymous):

-3x-9=-3x-3-5

OpenStudy (lilg132):

-3x - 9 = -3x - 3 -5

OpenStudy (lilg132):

bring the x's to one side and numbers on the other

OpenStudy (anonymous):

-3x-9=-3x-8

OpenStudy (anonymous):

9-8 is 1 isnt it?

OpenStudy (lilg132):

my mistake left the 3 out

OpenStudy (lilg132):

x = 9 - 8

OpenStudy (anonymous):

0x=1--->0=1 theres no solution

OpenStudy (anonymous):

thanks for the medal, lilg! :)